# When LLMs Play the Telephone Game: Cumulative Changes and Attractors in Iterated Cultural Transmissions

### v2

@@ 1 Introduction
Large language models (LLMs) are playing an increasingly significant role in the production of media content across various domains machine_culture. They are being used for academic writing buruk_academic_2023; agarwal_litllm_2024; cheng_have_2024; dergaa_human_2023; khraisha_can_2024; marlow_ghost_2021; cabanac_tortured_2021, journalism petridis_anglekindling_2023, story generation valentini_automatic_2023; xie_next_2023; zhao_more_2023, as chatbots on social media sadikoglu_evolution_2023 and in the workplace at large brynjolfsson_generative_2023; eloundou_gpts_2023. As LLMs become more performant, controllable, and more widespread, their impact on the creation and dissemination of information content humans consume is expected to grow even further machine_culture.
Given the implications of LLM usage for the production of cultural information, numerous researchers have studied the properties of LLM-generated content. In these studies, LLMs showed several biases with respect to gender acerbi2023large; salewski_-context_2024; wan_kelly_2023; haller_opiniongpt_2023; dong_disclosure_2024, race raj_breaking_2024, values which_humans; sivaprasad_exploring_2024, politics motoki_more_2023; agiza_analyzing_2024; haller_opiniongpt_2023, authority, fallacy oversight, and beauty chen_humans_2024. They were also found to generate texts at least as attractive marlow_ghost_2021 and compelling spitale_ai_2023 as human-generated ones and to display similar cognitive biases echterhoff_cognitive_2024.
@@
While single-turn behaviors of LLMs prompted with a human-generated prompt are under active scrutiny, little is known about the effect of iterated text interactions. Indeed, LLMs often use existing cultural content to generate new material, for example when writing scientific reviews khraisha_can_2024; agarwal_litllm_2024 or generating new stories based on existing examples xie_next_2023. As the share of AI-generated content increases, LLMs will be producing outputs using content that is already LLM-generated, making it crucial to study the consequences of this iterative process. Moreover, LLMs are increasingly being used in multi-agent settings de_zarza_emergent_2023; park_generative_2023; park_social_2022; xiao_simulating_2023; hua_war_2023; vezhnevets_generative_2023; chuang2023simulating and are already interacting with one another as chatbots on social media [^1] https://chirper.ai/.
Nonetheless, little is known about how populations of LLMs might self-organize, as other complex systems do. Research in complex systems traditionally studies how global-level patterns emerge from local interactions gleick_chaos_1988; mitchell_complexity_2009. Here, we ask whether multi-turn behaviors conditioned on LLM-generated content cause the appearance of new kinds of biases, undetectable in single-turn behaviors but accumulating across iterated interactions.
@@
To address this question, we take inspiration from the cultural evolution literature. The field of cultural evolution aims to provide causal explanations for the change of culture over time. In using the term culture, we borrow a definition traditionally used in the cultural evolution literature, where it refers to any type of socially transmitted information mesoudi_cultural_2016. This definition includes various types of cultural information (e.g. beliefs, skills, norms) and is independent of the medium of transmission (e.g. bodily imitation, oral or written transmission). Importantly, this way of defining culture is quite different from the view that is adopted by research works studying cultural values, or cultural alignment in LLMs, where culture is seen as an attribute of a specific human group (e.g. “Western culture”). This definition allows to use the same conceptual and methodological framework which was applied to study human derex_causal_2019 and non-human whiten_culture_2011 cultural evolution to study LLMs, which is why we adopt that definition for this work.
@@
In particular, we draw insights from a research tradition called cultural attraction theory (CAT) sperber_anthropology_1985; morin_how_2016; miton_cultural_2024. CAT aims to determine how non-random transformations of cultural information during transmission events may lead to the evolution of progressively more stable forms, referred to as attractors. Although the precise conceptualization of attractors varies across authors, an encompassing definition may be “theoretical posits that capture the way in which certain ideational variants are more likely to be the outcome of transformations than others.” buskell_what_2017. Importantly, the existence of attractors would imply that for a broad range of starting conditions, different variants of a given cultural trait will converge towards the same state, that is, to a cultural trait that possesses specific properties.
@@
Most experiments in CAT employ a transmission chain design mesoudi_experimental_2021, first introduced by bartlett_remembering_1932. In this setup, chains of participants receive, produce and transmit social information from and to each other in a sequential manner (as in the popular telephone game). This powerful and highly controlled design allows to evaluate the high-level patterns that emerge from the accumulation of directional changes during single-turn transmission events. Here, we adapt this design to study how culture evolves along chains of LLMs, rather than human participants (Figure 1).
@@
The code for reproducing the simulations, analyses and figures is available on our GitHub [^2] https://anonymous.4open.science/r/TelephoneGameLLM-D9E1.
@@
  • We introduce an experimental paradigm based on transmission chains to study the biases and attractors introduced by multi-turn LLM interactions (Section 3.1)
  • We introduce a method to statistically assess the additional effects of multi-turn interactions over single-turn ones. (Section 3.3)
  • We introduce a method to estimate the existence, position, and strength of potential cultural attractors (Section 3.4).
  • We study the evolution of text properties (positivity, difficulty, toxicity, and length) across iterated transmissions and uncover systematic biases and attractors that are specific to multi-turn interactions (Section 4.2)
  • We show how the LLM properties (model and size) and the task influence the position and strength of attractors, as well as the consequence of fine-tuning and temperature tuning (Section 4.3).
@@ 2 Related work
LLM-generated content is known to exhibit a variety of stereotypical biases dangers_parrots; weidinger2021ethical. In single-turn settings, LLMs perpetrate nadeem2020stereoset or even amplify human biases based on gender, nationality, race, and religion kotek2023gender. For instance, the GPT model was shown to exhibit cultural values similar to those of WEIRD (Western, Educated, Industrial, Rich, Democratic) cultures which_humans. LLMs trained through reinforcement learning with human feedback (RLHF) were found to overly express left-wing opinions on American politics — a tendency that, once formed, is difficult to avoid even after steering the model toward different demographic groups santurkar2023whose.
@@
Several studies have applied experimental designs used in cultural evolution to study knowledge and skill accumulation in groups of Reinforcement Learning agents cook_artificial_2024; schmitt_kickstarting_2018; open_ended_learning_team_open-ended_2021; prystawski_cultural_2023. Closer to the current study, populations of LLMs have also been studied machine_culture. Iterative chains of generative models trained on the preceding model’s output have been shown to sometimes collapse toward the most likely outputs while the tails of the original distribution disappear shumailov2023curse; peterson2024ai. The idea of using LLM-generated content to fine-tune the next generation has also been applied to groups of LLMs with various communication structures helm_tracking_2024. Similar to our approach, iterative chains with frozen (i.e., not re-trained) LLMs have been shown to express human-like biases in terms of gender stereotypes, positivity, and social, threat, and biology-related information acerbi2023large. Strong, but non-human-like biases for producing factual information have also been observed chuang2023simulating, stressing the importance of understanding the evolution of content in LLMs and the ways it might deviate from human cultural evolution.
@@
Parallel works (ren_language_2024; zhu_eliciting_2024) have studied similar questions using different methodologies, namely the framework of Bayesian Iterated Learning (BIL) (kirby_iterated_2014; griffiths_language_2007). Under this framework, attraction is conceptualized as convergence towards priors, whereas we conceptualize it as the fixed point of a recurrent sequence. Therefore, BIL allows to make prediction about how an agents’ hypothesis about a given phenomenon will evolve through social interaction, while our method captures how a specific cultural trait will evolve. Moreover, our framework enables us to avoid assuming that LLMs possess “priors” in the same way as humans do, which might be problematic (see kovac_large_2023; shanahan_role-play_2023).
@@ 3 Methods
Our telephone game experiments aim to study the possible attractors and biases that may accumulate across multiple turns of interactions between LLMs. This is done with a transmission chain design tracking the evolution LLM outputs as a function of the number of interactions in the LLM chain. This section introduces our transmission chain design (Section 3.1), the set of metrics used to study the evolution of text properties (Section 3.2) the effects of multi-turn interactions 3.3, and our method to characterize the properties of attractors (Section 3.4).
3.1 LLM transmission chains
In transmission chains, individual participants are ordered linearly. Each participant receives some information from the previous one, performs a task, and transmits new information to the next participant. Each agent is prompted with a task (instruction on how the text should be processed) and a text, which are concatenated and passed to the user message. The first agent is given a human-generated text and a task, and subsequent agents are given the same task and the text generated by the previous agent in the transmission chain:
$text_{i+1}=LLM(task,text_{i}),$ (1)
where text0 is the initial human-generated text and $LLM$ generates an output based on task $task$ and the previous agent’s text $x_{i}$. We run this process for 50 generations. Examples of texts evolving through generations are provided in Appendix A.
Initial texts ($text_{0}$)
We borrow human-generated text from various databases to provide the initial input to each transmission chain. Since we were interested in how variation in the initial text would impact the properties of the ensuing chain, human-generated texts spanned various types of content: scientific abstracts  [^3] https://huggingface.co/datasets/CCRss/arxiv_papers_cs, news articles  [^4] https://huggingface.co/datasets/RealTimeData/bbc_latest, and social media posts  [^5] https://huggingface.co/datasets/FredZhang7/toxi-text-3M/blob/e0e5b168b4a7e14e84f07271bfe1c6b42bc91ccd/multilingual-train-deduplicated.csv. As we are interested in the evolution of the toxicity, positivity, difficulty and length of generated texts, we sample the entire dataset to obtain a subset of 20 initial texts that covered the range of possible values for these properties. The exact method used to extract these texts is detailed in Appendix A.
Tasks
To determine the effects of instructions on the evolution of content over generations of LLMs, we prompt each chain of LLMs with three different tasks encompassing typical uses of LLMs:
@@ 3.1 LLM transmission chains
Tasks remained consistent within each chain. The exact prompt used for each task is reported in Appendix A.
Models
To assess whether and how cultural evolution dynamics are affected by the model specifications, we run identical experiments using five different models, all commonly used, from three different companies and with varying sizes: GPT-3.5-turbo-0125 (GPT3.5) , Llama3-8B-Instruct (Llama3-8B), Mistral-7B-Instruct-v0.2 (Mistral-7B), Llama3-70B-Instruct (Llama3-70B), Mixtral-8x7B-Instruct-v0.1 (Mixtral-8x7B). For inference, we used the OpenAI API (MIT License)  [^6] https://openai.com/index/openai-api/ to run GPT3.5 and the HuggingFace’s Transformer library wolf2019huggingface for other models (Apache Licence, v2.0). In our setup, transmission chains are always homogeneous with respect to the model, i.e. each chain is composed of a population of agents sharing the same underlying model.
@@ 3.2 Metrics
Text properties
Iterated transmissions may affect the generated text in several ways. We focus on four, orthogonal properties for each text which could be automatically measured (as opposed to requiring human annotators, which would have been impractical given the size of the output corpus):
  • Toxicity. Companies typically fine-tune LLMs to avoid the generation of toxic (i.e., dangerous or harmful) content. However, to our knowledge, this fine-tuning step focuses on single-turn dynamics, and the evolution of content with respect to its toxicity is currently understudied. We measure the toxicity of texts by quantifying the presence of rude, disrespectful, or unreasonable language, using a probability score that ranges from 0.0 (benign and non-toxic) to 1.0 (highly likely to be toxic), as estimated by the classifier introduced in Detoxify.
  • Positivity. Even when trained to avoid toxic content, LLMs have been shown to express similar positivity biases to humans, often favoring negative over positive information in preserving and generating new information acerbi2023large. To study whether positivity biases over transmission chains are affected by tasks and models, we measure the positivity of produced contents using the SentimentIntensityAnalyzer tool from NLTK hardeniya2016natural. It uses this information to calculate a sentiment score for the text, ranging from -1.0 (highly negative) to 1.0 (highly positive).
  • Difficulty. While LLMs are argued to benefit society by democratizing knowledge techstrongGenerativeNext, such positive outcomes are conditioned on the LLMs generating output that is accessible and inclusive to all kinds of audiences. However, whether text difficulty is preserved, increased, or reduced over transmission chains is currently unknown. We estimate text difficulty using the Gunning-Fog index bogert1985defense, which depends on the average sentence length and the percentage of difficult words. A standard interpretation of this index is that it estimates the years of formal education required to fully understand the text.
  • Length: A simple, yet crucial aspect of a piece of text is its length. As more and more content is generated by and from LLM outputs, cultural media may become populated with increasingly short (potentially incomplete) or long (potentially redundant, meaningless, or hard to process) material. We therefore assess the evolution of content length as measured by the character count of generated text.
We provide additional details about metrics in Appendix A.
@@ 3.3 Effect of multi-turn transmissions
One of our questions is how content evolves over multi-turn transmissions compared to single-turn settings. To address this point, we compare the distribution of a given property in the generated texts at the first generation to the distributions at subsequent generations. Thus, for each model and task, we look at the properties of each of the 100 generated texts (20 transmissions chains * 5 seeds) at each generation, which gives us a sample of 100 property values for each value. Using a Kolmogorov–Smirnov test ks_test, we then test whether the sample obtain at each generation comes from the same distribution as the sample obtained after the first generation. If we can confidently reject the hypothesis that the sample of property values at the end of the transmission chain comes from the same distribution as the sample obtained after the first generation, this would confirm that looking at outputs after a single-turn transmission is not enough for predicting output properties in a multi-turn setting.

### v3

# When LLMs Play the Telephone Game: Cultural Attractors as Conceptual Tools to Evaluate LLMs in Multi-turn Settings

@@ 1 Introduction
Large language models (LLMs) are playing an increasingly significant role in the production of media content across various domains (Brinkmann et al., 2023). They are being used for academic writing (Buruk, 2023; Agarwal et al., 2024; Cheng et al., 2024; Dergaa et al., 2023; Khraisha et al., 2024; Marlow and Wood, 2021; Cabanac et al., 2021), journalism (Petridis et al., 2023), story generation (Valentini et al., 2023; Xie et al., 2023; Zhao et al., 2023), as chatbots on social media (Sadikoğlu et al., 2023) and in the workplace at large (Brynjolfsson et al., 2023; Eloundou et al., 2023). As LLMs become more performant, controllable, and more widespread, their impact on the creation and dissemination of information content humans consume is expected to grow even further (Brinkmann et al., 2023).
Given the implications of LLM usage for the production of cultural information, numerous researchers have studied the properties of LLM-generated content. In these studies, LLMs showed several biases with respect to gender (Acerbi and Stubbersfield, 2023; Salewski et al., 2024; Wan et al., 2023; Haller et al., 2023; Dong et al., 2024), race (Raj et al., 2024), values (Atari et al., 2023; Sivaprasad et al., 2024), politics (Motoki et al., 2023; Agiza et al., 2024; Haller et al., 2023), authority, fallacy oversight, and beauty (Chen et al., 2024). They were also found to generate at least as attractive (Marlow and Wood, 2021) and compelling (Spitale et al., 2023) texts than humans and to display similar cognitive biases (Echterhoff et al., 2024).
@@
While single-turn behaviors of LLMs prompted with a human-generated prompt are under active scrutiny, little is known about the effect of iterated interactions. Indeed, LLMs often use existing cultural content to generate new ones, for example when writing scientific reviews (Khraisha et al., 2024; Agarwal et al., 2024) or generating new stories based on existing examples (Xie et al., 2023). As the share of AI-generated content increases, LLMs will be producing outputs using content that is already LLM-generated, making it crucial to study the consequence of this iterative process. Moreover, LLMs are increasingly being used in multi-agent settings (de Zarzà et al., 2023; Park et al., 2023; 2022; Xiao et al., 2023; Hua et al., 2023; Vezhnevets et al., 2023; Chuang et al., 2023) and are already interacting with one another as chatbots on social media [^1] https://chirper.ai/. Nonetheless, little is known about how populations of LLMs might self-organize, as other complex systems do. Research in complex systems traditionally studies how global-level patterns emerge from local interactions (Gleick and Hilborn, 1988; Mitchell, 2009). Here, we ask whether multi-turn behaviors conditioned on LLM-generated content cause the appearance of new kinds of biases, undetectable in single-turn behaviors but accumulating across iterated interactions.
@@
To address this question, we take inspiration from the cultural evolution literature. The field of cultural evolution aims to provide causal explanations for the change of culture over time. By cultural, we refer to the definition traditionally used in the cultural evolution literature, where it refers to any type of socially transmitted information (Mesoudi, 2016). This definition includes various types of cultural information (e.g. beliefs, skills, norms) and is independent of the medium of transmission (e.g. bodily imitation, oral or written transmission). Importantly, this way of defining culture is quite different from the view that is adopted by research works studying cultural values, or cultural alignment in LLMs, where culture is seen as an attribute of a specific human group (e.g. “Western Culture”). This definition allows to use the same conceptual and methodological framework which was applied to study human (Derex et al., 2019) and non-human (Whiten et al., 2011) cultural evolution to study LLMs, which is why we adopt that definition for this work.
@@
In particular, we draw insights from a research tradition called cultural attraction theory (CAT) (Sperber, 1985; Morin, 2016; Miton, 2024). CAT aims to determine how non-random transformations of cultural information during transmission events may lead to the evolution of progressively more stable forms, referred to as attractors. Although the precise conceptualization of attractors varies across authors, an encompassing definition may be "theoretical posits that capture the way in which certain ideational variants are more likely to be the outcome of transformations than others." (Buskell, 2017). Importantly, the existence of attractors would predict that for a broad range of starting conditions, different variants of a given cultural trait will converge towards the same attractor, that is, to a cultural trait that possesses specific properties.
@@
Most experiments in CAT employ a transmission chain design (Mesoudi, 2021), first introduced by (Bartlett, 1932). In this setup, chains of participants receive, produce and transmit social information from and to each other in a sequential manner (as in the popular telephone game). This powerful and highly controlled design allows to evaluate the high-level patterns that emerge from the accumulation of directional changes during single-turn transmission events. Here, we adapt this design to study how culture evolves along chains of LLMs, rather than human participants (Figure 1).
@@
The code for reproducing the simulations, analyses and figures is available on our GitHub [^2] https://github.com/jeremyperez2/TelephoneGameLLM
@@
  • We propose that there might be a gap in current LLM evaluations methods (single-turn evaluations might not be suited to assess the properties of multi-turn interactions)
  • We empirically confirm this hypothesis by showing that multi-turn interactions indeed often lead to distributions of text properties that are significantly different from what is observed after a single interaction (Section 4.2)
  • We introduce novel conceptual and methodological tools to fill this gap, grounded in research in cultural evolution, and in particular the concept of cultural attractor (Section 3.4).
  • We showcase the potential of this method by applying it to compare the effect of different tasks, of different models, of temperature, and of fine-tuning on the properties of multi-turn interactions. (Section 4.3).
  • We find several robust effects, such as the fact that less constrained tasks lead to stronger attractors, that some properties posses stronger attractors than others, and that fine-tuning can shift the position and modify the strength of attractors.
@@ 2 Related work
LLM-generated content is known to exhibit a variety of stereotypical biases (Bender et al., 2021; Weidinger et al., 2021; Bai et al., 2024). In single-turn settings, LLMs perpetrate (Nadeem et al., 2020) or even amplify human biases based on gender, nationality, race, and religion (Kotek et al., 2023). For instance, the GPT model was shown to exhibit cultural values similar to those of WEIRD (Western, Educated, Industrial, Rich, Democratic) cultures (Atari et al., 2023). LLMs trained through reinforcement learning with human feedback (RLHF) were found to overly express left-wing opinions on American politics — a tendency that, once formed, is difficult to avoid even after steering the model toward different demographic groups (Santurkar et al., 2023). Finally, a large body of work compared cognitive biases of models to those of humans (Coda-Forno et al. (2024); Liu et al. (2024); Binz and Schulz (2023)).
@@
Several studies have applied experimental designs used in cultural evolution to study knowledge and skill accumulation in groups of Reinforcement Learning agents (Cook et al., 2024; Schmitt et al., 2018; Team et al., 2021; Prystawski et al., 2023). Closer to the current study, populations of LLMs have also been studied (Brinkmann et al., 2023). Iterative chains of generative models trained on the preceding model’s output have been shown to sometimes collapse toward the most likely outputs while the tails of the original distribution disappear (Shumailov et al., 2024; Peterson, 2024; Gerstgrasser et al., 2024; Kazdan et al., 2024; Dohmatob et al., 2024). This idea of using LLM-generated content to fine-tune the next generation has also been applied to groups of LLMs with various communication structures (Helm et al., 2024). Similar to our approach, iterative chains with frozen (i.e., not re-trained) LLMs have been shown to express human-like biases in terms of gender stereotypes, positivity, and social, threat, and biology-related information (Acerbi and Stubbersfield, 2023). Strong, but non-human-like biases for producing factual information have also been observed (Chuang et al., 2023), stressing the importance of understanding the evolution of content in LLMs and the ways it might deviate from human cultural evolution.
@@
Parallel works ((Ren et al., 2024; Zhu and Griffiths, 2024)) have studied similar questions using different methodologies, namely the framework of Bayesian Iterated Learning (BIL) ((Kirby et al., 2014; Griffiths and Kalish, 2007)). Under this framework, attraction is conceptualized as convergence towards priors, whereas we conceptualize it as the fixed point of a recurrent sequence. Therefore, BIL allows to make prediction about how an agents’ hypothesis about a given phenomenon will evolve through social interaction, while our method captures how a specific cultural trait will evolve. Moreover, our framework enables us to avoid assuming that LLMs possess “priors” in the same way that humans do, which might be problematic (see Kovač et al. (2023); Shanahan et al. (2023)).
Evaluation of LLMs in social contexts
While not focusing on the effect of multi-turn interactions, an important body of work has started to study the social cognition of LLMs. These studies revealed that GPT-4 struggles to reach human performance with respect to social commonsense reasoning and strategic communication skills ((Zhou et al., 2024)). In a similar vein, Huang et al. (2024) found that manipulating the communication structure can make LLM collectives more resilient to the introduction of malicious agents.
@@ 3 Methods
3.1 Experimental paradigm: LLM transmission chains
In transmission chains, individual participants are ordered linearly. Each participant receives some information from the previous one, performs a task, and transmits new information to the next participant. Each agent is prompted with a task (instruction on how the text should be processed) and a text, which are concatenated and passed to the user message. The first agent is given a human-generated text and a task, and subsequent agents are given the same task and the text generated by the previous agent in the transmission chain: $text_{i+1}=LLM(task,text_{i})$, where text0 is the initial human-generated text and $LLM$ generates an output based on task $task$ and the previous agent’s text $x_{i}$. We run this process for 50 generations. Examples of texts evolving through generations are provided in Appendix Section A.
Initial texts ($text_{0}$) We borrow human-generated text from various databases to provide the initial input to each transmission chain. Since we were interested in how variation in the initial text would impact the properties of the ensuing chain, human-generated texts spanned various types of content: scientific abstracts As we are interested in the evolution of the toxicity, positivity, difficulty and length of generated texts, we sample the entire dataset to obtain a subset of 20 initial texts that covered the range of possible values for these properties. The exact method used to extract these texts is detailed in Appendix  A. As a robustness checks, we replicated our experiments with a larger set of 100 initial texts (Appendix B).
Tasks To determine the effects of instructions on the evolution of content over generations of LLMs, we prompt each chain of LLMs with three different tasks encompassing typical uses of LLMs:
@@ 3.1 Experimental paradigm: LLM transmission chains
Tasks remained consistent within each chain. The exact prompt used for each task is reported in Appendix Section A.
Models To assess whether and how cultural evolution dynamics are affected by the model specifications, we run identical experiments using six different models, all commonly used, from three different companies and with varying sizes: GPT-4o-mini, GPT-3.5-turbo-0125 (referred to as ”GPT3.5”) , Llama3-8B-Instruct (”Llama3-8B”), Mistral-7B-Instruct-v0.2 (”Mistral-7B”), Llama3-70B-Instruct (”Llama3-70B”), Mixtral-8x7B-Instruct-v0.1 (”Mixtral-8x7B”). For inference, we used the OpenAI API (The MIT License)  [^3] https://openai.com/index/openai-api/ to run GPT-4o-mini and GPT3.5 and the HuggingFace’s Transformer library (Wolf et al., 2019) for other models (Apache Licence, v2.0).
@@ 3.2 Metrics
Iterated transmissions may affect the generated text in several ways. We focus on four, orthogonal properties for each text which could be automatically measured:
Toxicity. Companies typically fine-tune LLMs to avoid the generation of toxic (i.e., dangerous or harmful) content. However, to our knowledge, this fine-tuning step focuses on single-turn dynamics, and the evolution of content with respect to its toxicity is currently understudied. We measure the toxicity of texts by quantifying the presence of rude, disrespectful, or unreasonable language, using a probability score that ranges from 0.0 (benign and non-toxic) to 1.0 (highly likely to be toxic), as estimated by the classifier introduced in Hanu and Unitary team (2020).
Positivity. Even when trained to avoid toxic content, LLMs have been shown to express similar negativity biases to humans, often favoring negative over positive information in preserving and generating new information (Acerbi and Stubbersfield, 2023). To study whether positivity biases over transmission chains are affected by tasks and models, we measure the positivity of produced contents using the SentimentIntensityAnalyzer tool from NLTK (Hardeniya et al., 2016). It uses this information to calculate a sentiment score for the text, ranging from -1.0 (highly negative) to 1.0 (highly positive).
Difficulty. While LLMs are argued to benefit society by democratizing knowledge (Weiss, ), such positive outcomes are conditioned on the LLMs generating output that is accessible and inclusive to all kinds of audiences. However, whether text difficulty is preserved, increased, or reduced over transmission chains is currently unknown. We estimate text difficulty using the Gunning-Fog index (Bogert, 1985), which depends on the average sentence length and the percentage of difficult words. A standard interpretation of this index is that it estimates the years of formal education required to fully understand the text.
Length: A simple, yet crucial aspect of a piece of text is its length. As more and more content is generated by and from LLM outputs, cultural media may become populated with increasingly short (potentially incomplete) or long (potentially redundant, meaningless, or hard to process) material. We therefore assess the evolution of content length as measured by the character count.
We provide additional details about metrics in Appendix Section A.
@@ 3.3 Effect of multi-turn transmissions
One of our questions is how content evolves over multi-turn transmissions compared to single-turn settings. To address this point, we compare the distribution of a given property in the generated texts at the first generation to the distributions at subsequent generations. Thus for each model and task, we look at the properties of each of the 100 generated texts (20 transmissions chains * 5 seeds) at each generation, which gives us a sample of 100 property values for each value. Using a Kolmogorov–Smirnov test (Massey, 1951), we then test whether the sample obtain at each generation comes from the same distribution as the sample obtained after the first generation. If we can confidently reject the hypothesis that the sample of property values at the end of the transmission chain comes from the same distribution as the sample obtained after the first generation, this would confirm that looking at outputs after a single-turn transmission is not enough for predicting output properties in a multi-turn setting.